OpenStudy (anonymous):

what is the surface area of a conical grain storage tank that has a height of 37 meters and a diameter of 16 meters round the answer to the nearest square meter A. 2,831 square meters B. 2,664 square meters C. 1,152 square meters D. 1,131 square meters

OpenStudy (anonymous):

im still confused @alekos

OpenStudy (anonymous):

okay so the base of this grain storage is a circle right? since its conical so just use the circle area formula: pi*r*^2

OpenStudy (anonymous):

that would be the area of the base then multiply by the height of 37 meters

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

im still confused would h be 37 and r is 16???????

OpenStudy (anonymous):

i got 2368pi/3

OpenStudy (anonymous):

is that right?

OpenStudy (alekos):

stand by, they want surface area not volume

OpenStudy (alekos):

Surface area of a cone \[pir(r+\sqrt{h ^{2}+r^{2}})\]

OpenStudy (anonymous):

I have to go for like 10 to 15 mins but i will be back

OpenStudy (alekos):

r=8 and h=37 so we have \[8\pi(8+\sqrt{37^{2}+8^{2}}) \]

OpenStudy (anonymous):

this time i got 1,152

OpenStudy (anonymous):

diameter is 16 and the radius is half of the diameter so the radius is 8

OpenStudy (alekos):

That is correct